# Dolphin v1.0 Technical Report

### v2

@@ Abstract
Ultrasound is vital in modern medicine but faces challenges like operator dependence, image noise, and dynamic real-time scanning, which hinder effective AI integration. Though large multimodal models excel in other medical imaging areas, they struggle with ultrasound’s unique complexities. To bridge this gap, we introduce Dolphin v1.0 (V1) and its reasoning-augmented version, Dolphin R1—the first large-scale multimodal ultrasound foundation models designed to unify diverse clinical tasks within a single vision-language framework. To tackle the variability, noise, and operator dependence inherent in ultrasound imaging, we curated a groundbreaking multimodal dataset at the 2-million-scale, integrating textbook knowledge, public ultrasound data, synthetic knowledge-distilled samples, and general multimodal corpora. This comprehensive dataset ensures the models achieve robust perception, strong generalization, and broad clinical adaptability across medical domains.
The Dolphin series adopts a progressive three-stage training strategy that combines domain-specialized pretraining, instruction-driven alignment, and reinforcement-based refinement. The baseline Dolphin v1.0 establishes reliable performance in classification, detection, regression, and report generation, offering a strong general-purpose model for clinical ultrasound analysis. Building upon this foundation, Dolphin R1 leverages reinforcement learning with verifiable ultrasound-specific reward signals, thereby enabling deeper diagnostic inference, enhanced reasoning transparency, and more interpretable decision pathways.
To systematically evaluate performance,We adopt U2-Bench for evaluation eight representative ultrasound tasks, ranging from lesion localization and organ detection to clinical value estimation and structured report generation. Experimental results demonstrate that Dolphin R1 achieves a U2-score of 0.5835—more than twice the score of the second-best model (0.2968)—establishing a new state of the art in multimodal ultrasound understanding. Meanwhile, Dolphin v1.0 also delivers competitive performance, validating the effectiveness of the unified training framework. Further comparison between standard mode (V1) and deep reasoning mode (R1) reveals that reasoning-enhanced training significantly boosts diagnostic accuracy, consistency, and interpretability. These findings underscore the importance of integrating reasoning into foundation models for high-stakes medical domains, providing a scalable paradigm for future developments in intelligent ultrasound analysis.
@@ Introduction
Ultrasound (US) is one of the most widely utilized imaging modalities in clinical practice, playing a critical role in obstetrics, cardiology, emergency medicine, and resource-limited settings due to its real-time imaging capability, portability, and cost-effectiveness(Hewson & Bedforth (2023)). However, despite its broad applicability, ultrasound interpretation remains highly challenging compared to other medical imaging modalities such as computed tomography (CT), magnetic resonance imaging (MRI), and positron emission tomography (PET). These modalities provide high spatial resolution, standardized anatomical views, and consistent image quality, whereas ultrasound is inherently operator-dependent, susceptible to imaging artifacts, and exhibits significant variability in acquisition quality(Sharma et al. (2021)).
@@
Currently, multimodal large models have emerged as a powerful paradigm in artificial intelligence research(OpenAI (2023), Bai et al. (2025), Lin et al. (2023), Wu et al. (2024)), demonstrating remarkable capabilities in processing and integrating information across diverse modalities such as text, images, audio, and video. Currently, Multimodal large models have demonstrated significant potential in the medical field, offering advanced capabilities for integrating and analyzing diverse data types, including medical images (MRI,CT,PET; Lee et al. (2024); Volkov et al. (2025); Shui et al. (2025)), electronic health records(AlSaad et al. (2024)), genomic data(Fallahpour et al. (2025)), and clinical notes(Han et al. (2024)). While these models have achieved remarkable success in processing static, high-resolution medical images and structured clinical data, their application to ultrasound imaging presents unique challenges that remain largely unaddressed.
@@
The translation of multimodal approaches to ultrasound is hindered by several critical limitations. First, unlike the standardized acquisition protocols for CT or MRI, ultrasound examinations are highly operator-dependent, resulting in significant variability in image quality and anatomical views (Stasi & Ruoti (2015)). Second, the dynamic nature of ultrasound video sequences - often containing crucial temporal information about tissue motion and blood flow - is fundamentally different from the static images that current multimodal models typically process (Szabo (2013)). Third, the characteristic speckle noise and lower signal-to-noise ratio of ultrasound images compared to other modalities create additional challenges for feature extraction and interpretation (Loizou & Pattichis (2022)). Furthermore, existing multimodal models struggle with the sparse and often subjective nature of ultrasound reports, which typically lack the detailed anatomical descriptions found in radiology reports for other imaging modalities(Wang et al. (2024)). Given the aforementioned challenges, developing large multimodal models specifically tailored for ultrasound imaging has become an urgent necessity. In recent years, several studies have attempted to apply ultrasound images in the research of large-scale multimodal models. The EchoCLIP model proposed by Christensen et al. (2024) learns the correspondence between cardiac ultrasound images and expert-annotated cardiac case reports, thereby achieving an understanding of cardiac-related semantic information. Subsequently, Guo et al. (2024) proposed LLaVA-Ultra, which integrates the CLIP and SAM encoders and constructs a general-purpose large model for the ultrasound domain through domain-specific fine-tuning. It has outperformed LLaVA-Med(Li et al. (2023)) on benchmark tests across various general medical tasks.
Although multimodal large models have been increasingly explored in the ultrasound domain, existing approaches often lack comprehensive generalization and in-depth medical reasoning capabilities. To address these challenges, we introduce Dolphin V1 and Dolphin R1, the world’s first general-purpose multimodal ultrasound large models. Dolphin V1 is a vision-language model capable of jointly understanding ultrasound images, videos, dynamic scanning sequences, and associated clinical text, while Dolphin R1 further enhances the model’s reasoning and diagnostic inference capabilities.
@@
To support model development, we curated a large-scale multimodal ultrasound dataset containing over 2 million high-quality samples sourced from medical textbooks, public ultrasound repositories, general multimodal data, and knowledge-distilled synthetic data. This diverse and extensive dataset underpins the models’ robustness and clinical adaptability. Additionally, we propose the Dolphin Ultrasound Data Protocol (DUDP), a standardized data interface that unifies and formalizes a wide range of ultrasound tasks, laying the groundwork for future research and development in this domain.
In our work, we first unify a wide range of complex ultrasound tasks by abstracting them into four core task categories: classification, detection, regression, and generation. To effectively train the model, we employ a progressive three-stage training strategy, which simultaneously emphasizes the model’s generalization capability and medical understanding. Furthermore, we incorporate large-scale reinforcement learning based on ultrasound data, endowing the model with advanced reasoning abilities for complex diagnostic scenarios.
The experimental results indicate that our model outperforms all compared methods across the eight diverse ultrasound tasks defined in the U2-Bench benchmark. It achieves superior performance on a wide range of tasks, including disease diagnosis, organ detection, and medical report generation, demonstrating robustness and generalization capability in both low-level perception and high-level reasoning. Notably, our model achieves a U2-Bench score of 0.5835, significantly surpassing the second-best performing model, Gemini-2.5 Pro, which scores 0.2968. This result establishes a new state of the art in ultrasound-based multimodal modeling.
@@
  • We construct a large-scale, multimodal ultrasound dataset comprising over 2 million high-quality samples, covering both supervised fine-tuning (SFT) and reinforcement learning (RL) data.
  • We propose a progressive three-stage training framework that effectively enhances the model’s generalization capability and medical reasoning.
  • Our model achieves state-of-the-art performance on the U2-Bench benchmark, significantly outperforming existing methods across a wide range of ultrasound tasks, demonstrating strong low-level perception and high-level reasoning abilities.
These advancements collectively pave the way for significant progress in ultrasound-based medical AI, fostering more accurate, efficient, and intelligent clinical decision-making.
@@ 2.1 Post-training data
To enable the model to possess expert-level ultrasound image recognition capabilities, we constructed a large-scale, high-quality ultrasound image dataset. Specifically, our data is primarily sourced from the following four aspects: (1) Ultrasound textbooks and clinical guidelines - to enhance the model’s fundamental ultrasound recognition capabilities; (2) Publicly available ultrasound datasets covering classification, segmentation, detection, measurement, and other tasks - to improve the model’s disease diagnostic abilities; (3) General medical data - to enhance the model’s generalization and conversational capabilities; (4) Distilled knowledge data - to strengthen the model’s instruction-following and formatted output abilities.
During this process, we also constructed synthetic ultrasound data through the following approaches: (1) Question templates – Generated deterministic responses for structured data construction. (2) VQA data – Created using GPT-generated plausible options for ultrasound visual question answering. (3) Knowledge distillation data – Derived ultrasound datasets through model-based knowledge transfer, following the methodology of Xu et al. Xu et al. (2023a) on Evol-Instruct to iteratively enhance the quality of distilled data.
@@
Our data generation pipeline is meticulously designed to minimize the risk of hallucination and ensure high fidelity in clinical representation. The process begins with a well-defined task framework and template construction, which is followed by an initial review by board-certified physicians to validate the clinical relevance and diagnostic coherence of the proposed structure. Upon achieving expert consensus, synthetic data is generated in accordance with the predefined templates. This synthesized data then undergoes a second round of rigorous evaluation by medical professionals, who provide detailed feedback to guide iterative refinement and correction. Only after substantial alignment with clinical standards is achieved does the data proceed to a final validation stage, where it is independently assessed by physicians for inclusion in the training corpus. In total, this systematic approach incorporates three iterative rounds of cross-validation, ensuring robust clinical plausibility, diagnostic accuracy, and methodological rigor throughout the data creation process.
@@
Figure 1: Distribution of the Dolphin Ultrasound Large-Scale Dataset
2.1.1 Ultrasound teaching material data
@@
We collected a large number of classic ultrasound textbooks and industry guidelines, uniformly converted these materials into Markdown format, manually filtered out non-substantive content such as covers, back covers, and margins, and corrected grammatical errors and formulas that were inaccurately converted. Based on these ultrasound textbooks, we constructed the following three datasets: the Ultrasound Question-Answering Dataset, the Ultrasound Caption Dataset, and the Ultrasound Image-Text Explanation Dataset.
@@
Ultrasound Question-Answering Dataset: We segmented the textbook data into chunks and employed GPT to generate contextually appropriate questions for each segment. Through self-distillation, we regenerated corresponding answers to ensure strict question-answer consistency.
@@
Ultrasound Caption Dataset: The ultrasound textbook corpus contains systematic image-caption pairs, which we processed into a structured dataset through automated format matching to ensure precise image-annotation correspondence. It should be noted that our constructed caption dataset contains cases with multiple images, which helps enhance the model’s multi-image comprehension capability.
@@
Ultrasound Image-Text Explanation Dataset: The ultrasound textbooks contain detailed textual descriptions of medical images. We first identify the text passages corresponding to each image, then process these descriptions through GPT using self-distillation to generate high-quality question-answer pairs. This approach ensures data accuracy while simultaneously enhancing the model’s response diversity and instruction-following capabilities.
2.1.2 Public ultrasound datasets
@@
To enhance the model’s diagnostic capabilities across various tasks, we collected a substantial number of publicly available ultrasound datasets and established a unified data processing pipeline. We developed a unified file standard called Ultrasound Data Protocol (UDP) to harmonize diverse ultrasound tasks by converting them into a standardized JSON format. The protocol comprehensively handles diverse ultrasound tasks, supporting multimodal data types including ultrasound images, video sequences, segmentation masks, and associated quantitative measurements.
@@
Specifically, we categorize ultrasound tasks into the following standardized classes: (1) Anatomical Recognition (2) Standard View Identification (3) Diagnostic Classification (4) Tissue Segmentation (5) Biometric Measurement. Corresponding ultrasound question-answering datasets were constructed for each task category. To mitigate hallucination risks, we developed over a hundred fixed question templates for each task category. For each sample, we randomly selected a question template and populated it with JSON-formatted content, thereby constructing a high-quality ultrasound question-answering dataset. Figure 2 presents an authentic diagnostic data sample from our constructed dataset.
@@
Furthermore, to ensure data diversity, we employed GPT for partial data augmentation and utilized another LLM to evaluate response quality, guaranteeing that the JSON-formatted content remains embedded within the generated responses.
@@
Figure 2: Overview of the Dolphin Ultrasound Public Data Pipeline. The framework includes data collection, expert-guided filtering, and standardization via the Dolphin Ultrasound Data Protocol (DUDP), ensuring consistency across multimodal sources. Tasks are unified into four categories—classification, detection, regression, and generation—covering diverse clinical scenarios. The dataset distribution demonstrates balanced coverage across major anatomical regions, supporting robust model generalization.
2.1.3 General data
To enhance the model’s general capabilities while incorporating ultrasound-specific knowledge, we selected corresponding public datasets based on different competency requirements. The selected public datasets are primarily utilized to enhance the following model capabilities: (1) Comprehension; (2) Conversational; (3) Deep reasoning; (4) Medical; (5) Image-text dialogue; The primary public datasets adopted in our study are as follows: LLaVA-One-Vision. (Li et al., 2024), AutoIF. (Dong et al., 2024), Alpaca-GPT4 (Liu et al., 2023), Medical-O1-Reasoning (Chen et al., 2024). These datasets have been meticulously curated to comprehensively enhance the model’s performance across multiple dimensions.
2.1.4 Distilled knowledge data

### v3

# Dolphin Technical Report

@@ Abstract
Ultrasound is vital in modern medicine but faces challenges like operator dependence, image noise, and dynamic real-time scanning, which hinder effective AI integration. Though large multimodal models excel in other medical imaging areas, they struggle with ultrasound’s unique complexities. To bridge this gap, we introduce Dolphin V1 and its reasoning-augmented version, Dolphin R1— the first multimodal large language models (MLLMs) dedicated to ultrasound, designed to unify diverse clinical tasks within a single vision-language framework. We curate a multimodal dataset at the 2-million-scale, integrating textbook knowledge, diagnostic ultrasound data, synthetic knowledge-distilled samples, and general multimodal corpora. This comprehensive dataset ensures the models achieve robust perception, strong generalization, and broad clinical adaptability across medical domains.
The Dolphin series adopts a progressive three-stage training strategy that combines domain-specialized pretraining, instruction-driven alignment, and reinforcement-based refinement. The baseline Dolphin V1 establishes reliable performance in classification, detection, regression, and text generation, offering a strong general-purpose model for clinical ultrasound analysis. Building upon this foundation, Dolphin R1 leverages reinforcement learning with verifiable ultrasound-specific reward signals, thereby enabling deeper diagnostic inference, enhanced reasoning transparency, and more interpretable decision pathways.
To evaluate the models, we employ U2-Bench for ultrasound understanding and additionally assess performance on other medical benchmarks and general benchmarks. Experimental results demonstrate that Dolphin R1 achieves a U2-Score of 0.5835, more than twice the score of the second-best model (0.2968), establishing a new state of the art in multimodal ultrasound understanding. On broader benchmarks, Dolphin also achieves competitive or superior performance, with strong gains in general perception abilities (MME) and medical temporal reasoning (MedFrameQA). Further comparison between standard mode (V1) and deep reasoning mode (R1) reveals that reasoning-enhanced training significantly boosts diagnostic accuracy, consistency, and interpretability. These findings underscore the importance of integrating reasoning into MLLMs for medical domains. Together, we provide a scalable paradigm for future developments in intelligent ultrasound analysis.
@@ Introduction
Currently, multimodal large language models (MLLMs) have demonstrated significant potential in the medical field, offering advanced capabilities for integrating and analyzing diverse data types, including medical images (MRI,CT,PET; Lee et al. (2024); Volkov et al. (2025); Shui et al. (2025)), electronic health records (AlSaad et al. (2024)), genomic data (Fallahpour et al. (2025)), and clinical notes (Han et al. (2024)). While these models have achieved remarkable success in processing static, high-resolution medical images and structured clinical data, their application to ultrasound imaging presents unique challenges that remain largely unaddressed.
@@
Ultrasound is widely used in clinical practice due to its real-time capability, portability, safety, and cost-effectiveness (Hewson & Bedforth, 2023). However, the translation of multimodal approaches to ultrasound is hindered by several critical limitations. First, unlike the standardized acquisition protocols for CT or MRI, ultrasound examinations are highly operator-dependent, resulting in significant variability in image quality and anatomical views (Stasi & Ruoti (2015)). Second, the dynamic nature of ultrasound video sequences - often containing crucial temporal information about tissue motion and blood flow - is fundamentally different from the static images that current MLLMs typically process (Szabo (2013)). Third, the characteristic speckle noise and lower signal-to-noise ratio of ultrasound images compared to other modalities create additional challenges for feature extraction and interpretation (Loizou & Pattichis (2022)). Furthermore, existing MLLMs struggle with the sparse and often subjective nature of ultrasound reports, which typically lack the detailed anatomical descriptions found in radiology reports for other imaging modalities (Wang et al. (2024)). These challenges highlight the need for MLLMs capable of processing the unique characteristics of ultrasound imaging.
Medical ultrasound understanding requires precise perception of visual features, along with sophisticated reasoning to interpret dynamic, operator-dependent imaging and derive clinically meaningful conclusions. However, most existing multimodal medical models treat ultrasound tasks in isolation and focus primarily on perceptual accuracy, neglecting the development of systematic reasoning capabilities. Moreover, the scarcity of real-world diagnostic reasoning data in clinical settings poses a significant barrier to training models with robust medical reasoning skills. Recognizing these challenges, we introduce Dolphin, a general-purpose multimodal ultrasound understanding model capable of jointly interpreting ultrasound images, videos, dynamic scanning sequences, and associated clinical text. By leveraging a three-stage training framework with hybrid reasoning, Dolphin achieves deep cognitive capabilities in ultrasound analysis.
@@
To support model development, we curate a large-scale multimodal ultrasound dataset containing over 2 million high-quality samples sourced from medical textbooks, public repositories, clinical guidelines, and synthetic data. Additionally, we propose the Dolphin Ultrasound Data Protocol (DUDP), a standardized data interface that unifies and formalizes a wide range of ultrasound tasks, unifying ultrasound tasks into four categories: classification, detection, regression, and generation.
To effectively train a medical ultrasound language model with strong generalization and diagnostic reasoning capabilities, we adopt a progressive three-stage training strategy. The first stage leverages post-training on mixture of ultrasound-specific and general-domain data to build both medical foundation and broad language understanding. The second stage focuses on instruction tuning with multi-turn dialogues and deep-reasoning tasks, enhancing the model’s ability to engage in complex clinical discussions and perform step-by-step analysis. The final stage introduces UARPO, a reinforcement learning method guided by Ultrasound Answer Reward (UAR), which combines assessment of formatted answers and alignment with true diagnostic outcomes.
Dolphin achieves state-of-the-art performance with a U2-Bench (Le et al., 2025) score of 0.5835, outperforming all compared methods across eight ultrasound tasks. During the first stage of post-training, we train the model using easily accessible ultrasound QA data along with pure-text deep reasoning datasets from other domains. We observe that the model spontaneously developed a deep reasoning capability in ultrasound vision. We unify vanilla and deep reasoning within a Bayesian framework by modeling the latent chain-of-thought variable $c$. By introducing mode-dependent priors $P_{\theta}(c\mid x,m)$, we theoretically characterize cross-domain reasoning transfer: although deep reasoning data lacks target-domain samples, structured reasoning patterns generalize via shared parameters and prior bias. Our results highlight a critical threshold in reasoning data volume: effective reasoning emerges only when training exceeds this point. Meanwhile, experimental results show that deep reasoning mode improves diagnostic accuracy by 2.4%, reduces measurement RMSE by 10.6%, and increases detection accuracy by 16%. We find that base model performance determines the UARPO phase ceiling, with stronger foundation models achieving higher reinforcement learning gains.
@@
  • We introduce Dolphin, the first multimodal large language model tailored for ultrasound reasoning, with chat and reasoning versions. We also build a large-scale ultrasound dataset of over 2 million samples covering 16 body regions to enhance the model’s foundational capabilities.
  • We adopt a three-stage training strategy with mixed reasoning and UARPO-based reinforcement learning, using deep-reasoning data to improve the model’s visual reasoning for ultrasound. Ultimately, Dolphin successfully develops an emergent capacity for deep reasoning about ultrasound.
  • Dolphin achieves state-of-the-art performance on U2-Bench, surpassing all existing models across all six tasks with a remarkable U2-Score of 0.58, while also demonstrating strong generalization capabilities on mainstream benchmark datasets. Notably, Dolphin has been successfully deployed on real-world ultrasound devices, marking a significant step toward practical clinical integration by bringing AI directly into physicians’ daily workflows.
@@ 2.1 Post-training data
To enable the model to possess expert-level ultrasound image recognition capabilities, we constructed a large-scale, high-quality ultrasound multimodal dataset. Specifically, our data is primarily covering the following four aspects: (1) Ultrasound Knowledge-Oriented Data - to enhance the model’s fundamental knowledge about ultrasound; (2) Ultrasound Diagnostic-Oriented Data - covering classification, detection, regression and other tasks, to improve the model’s disease diagnostic abilities; (3) Ultrasound Domain Distillation Data - to strengthen the model’s instruction-following and formatted output abilities; (4) General Purpose Data - to enhance the model’s general and conversational capabilities. An illustration is included in Figure 1.
@@
Our data generation pipeline was meticulously designed to minimize the risk of hallucination and ensure high fidelity in clinical representation. The process began with a well-defined task framework and template construction, which was followed by an initial review by board-certified physicians to validate the clinical relevance and diagnostic coherence of the proposed structure. Upon achieving expert consensus, synthetic data was generated in accordance with the predefined templates. This synthesized data then underwent a second round of rigorous evaluation by medical professionals, who provided detailed feedback to guide iterative refinement and correction. After achieving substantial alignment with clinical standards, the data advanced to a final validation stage, where it was independently reviewed by physicians before inclusion in the training corpus. In total, this systematic approach incorporated three iterative rounds of cross-validation, ensuring robust clinical plausibility, diagnostic accuracy, and methodological rigor throughout the data creation process.
@@
Figure 1: Data types of the Dolphin Ultrasound Large-Scale Dataset
2.1.1 Ultrasound Knowledge-Oriented Data
@@
We collected a large number of classic ultrasound textbooks and industry guidelines, uniformly converted these materials into Markdown format, manually filtered out non-substantive content such as covers, back covers, and margins, and corrected grammatical errors and formulas that were inaccurately converted. Based on these ultrasound textbooks, we constructed the following three datasets: the Ultrasound Question Answering, the Ultrasound Caption, and the Ultrasound Image-Text Explanation.
@@
Ultrasound Question Answering. We segmented the textbook data into chunks and employed GPT to generate contextually appropriate questions for each segment. Through self-distillation, we regenerated corresponding answers to ensure strict question-answer consistency.
@@
Ultrasound Caption. The ultrasound textbook corpus contains systematic image-caption pairs, which we processed into a structured dataset through automated format matching to ensure precise image-annotation correspondence. It should be noted that our constructed caption dataset contains cases with multiple images, which helps enhance the model’s multi-image comprehension capability.
@@
Ultrasound Image-Text Explanation. The ultrasound textbooks contained detailed textual descriptions of medical images. We first identified the text passages corresponding to each image, then processed these descriptions through GPT using self-distillation to generate high-quality question-answer pairs. This approach ensures data accuracy while simultaneously enhancing the model’s response diversity and instruction-following capabilities.
2.1.2 Ultrasound Diagnostic-Oriented Data
@@
To enhance the model’s diagnostic capabilities across various tasks, we collected a substantial number of diagnostic-oriented ultrasound datasets and established a unified data processing pipeline. We developed a unified file standard called Dolphin Ultrasound Data Protocol (DUDP) to harmonize diverse ultrasound tasks by converting them into a standardized JSON format. The protocol comprehensively handles diverse ultrasound tasks, supporting multimodal data types including ultrasound images, video sequences, segmentation masks, and associated quantitative measurements.
@@
Specifically, we categorized ultrasound tasks into the following standardized classes: (1) Anatomical Recognition (2) Standard View Identification (3) Diagnostic Classification (4) Tissue Segmentation (5) Biometric Measurement. Corresponding ultrasound question-answering datasets were constructed for each task category. To mitigate hallucination risks, we developed over a hundred fixed question templates for each task category. For each sample, we randomly selected a question template and populated it with JSON-formatted content, thereby constructing a high-quality ultrasound question-answering dataset. Figure 2 presents an authentic diagnostic data sample from our constructed dataset.
@@
Furthermore, to ensure data diversity, we employed GPT for partial data augmentation and leveraged an LLM-based rating scheme to evaluate response quality, guaranteeing that the JSON-formatted content remains embedded within the generated responses.
@@
Figure 2: Overview of the Dolphin Ultrasound Diagnostic Data Pipeline. The framework includes data collection, expert-guided filtering, and standardization via the Dolphin Ultrasound Data Protocol (DUDP), ensuring consistency across multimodal sources. Tasks are unified into four categories—classification, detection, regression, and generation—covering diverse clinical scenarios. The dataset distribution demonstrates balanced coverage across major anatomical regions, supporting robust model generalization.
2.1.3 Ultrasound Domain Distillation Data